OpenStudy (anonymous):

What would be the slope and y intercept of 6x+7y=58?

OpenStudy (anonymous):

you need to isolate y on one side, and get everything else on the other side

OpenStudy (anonymous):

y=-6/7x + 8 is what I got.

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

get your equation in the form y = mx+b. b is the y intercept and m is the slope

OpenStudy (anonymous):

58/7 is not 8

OpenStudy (anonymous):

When i divide it, it comes out 8.3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you cannot round off your values

OpenStudy (anonymous):

Would I keep it as a decimal?

OpenStudy (anonymous):

fraction

OpenStudy (anonymous):

it is up to you. generally, it is more elegant to keep as fraction

OpenStudy (anonymous):

So, -6/7 would be the slope and 58/7 would be the y intercept?

OpenStudy (anonymous):

That is a very true statement

OpenStudy (anonymous):

indeed it would

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

you are welcome